In a linear double stranded DNA 6.2 kilo base pairs are present. If thymine is 2160 then give the no. Of guanine base 1) 8400 2)4040 3)4100 4)3860

Grade:12

2 Answers

AYUSHI ROYCHOWDHURY
6 years ago
Hiiiii I am AYUSHI ROYCHOWDHURY; I am here to give you the answer.
6.2kilo base pair=6.2 multiply with 1000= 6200 base pair
T+G=6200
2160+G=6200
G= 6200-2160
G= 4040 base pair
Simran Vinaik
6 years ago
Double stranded DNA has 6.2 kilo base pairs. so it will have 6200 base pairs or 12400 bases. Out of these 12400 bases, 2160 are Thymine.
According to Chargaff’s rule in double stranded DNA,
  1. number of Thymine will be equal to Adenine.
  2. number of Guanine wil be equal to Cytosine
A+T= 2160+2160= 4320

so G+C= 12400-4320
= 8080

thus no of guanine bases will be 8080/ 2= 4040